what is the inverse function of
*e= refers to the number
y= 1+ e^x / 1- e^x

Answer by venugopalramana(3286)

what is the inverse function of
*e= refers to the number
y= 1+ e^x / 1- e^x
(Y-1)/1 = E^X/(1-E^X)
(Y-1)/(Y-1+1)=E^X/(E^X+1-E^X)
(Y-1)/Y=E^X
X= LN{(Y-1)/Y}
HENCE INVERSE FUNCION IS
LN{(X-1)/X}